OpenStudy (anonymous):

Suppose you are drawing the doghouse and the scale you want to use is 1 in : 3 ft. What should the width of the doghouse be on the drawing, if the actual doghouse is 2 ft. 3 in. wide? A.) 0.75 in. OR B.) 2.3 in.

OpenStudy (anonymous):

a

OpenStudy (anonymous):

you know your scale is 1in - 3 ft, so you know a 2'3" doghouse is going to be smaller than an inch on the drawling.

OpenStudy (anonymous):

u can also do it by ratios: 1/3 = x / 2.25 x = 2.25 /3 = 0.75

OpenStudy (anonymous):

A